Steve’s Helicopter Rides takes tourists on helicopter tours of Hawaii. The price of each ticket is $150 and the variable costs average $60 per person. Steve’s monthly fixed costs are $58,500.

  1. If annual sales increase by $100,000, by what amount will operating income increase?
  2. If annual sales decrease by $50,000, by what amount will operating income decrease?
Answer #1

Contribution margin ratio = 150-60/150 = 60%

a) If annual Sales increase by $100000, operating income increase by (100000*60%) = 60000

b) If annual sales decrease by $50000 operating income decrease by (50000*60%) = -30000